Two vectors a and b have magnitudes 8 units and 12 units respectively their resultant is perpendicular to a find the angle between two vectors

Grade:12

1 Answers

ManpreetK
6 years ago
α=angle between resultant and vector of 8 magnitude. tanα=Bsinβ \ A+Bcosβ tan 90 = 12 sinβ \8 + 12cosβ1/tan90 =8+12cos β \8sinβ cot90 = 0So, 8 +12cos β=0 cos β = -8/12β= 131.81